Night Report 2017-01-03

Observers: Ken Smith, Aleksandar Cikota

TAT on duty: Janet Chen, Joe Lyman

Support team: QUB, TBD, Mark Magee

Daytime calibrations:


Weather: clear, 26% humidity, but strong wind, ~16m/s. Can't open the dome because of wind restriction.

02:06 UT: image analysis

02:22 UT: SOFI follow up of SN2016ije. Wind ~14 m/s.

03:35 UT: switch to EFOSC. Wind ~13.8 m/s. Seeing 1.37". 14% humidity. Pointing restrictions.

03:45 UT: image analysis

03:55 UT: Problem with telescope. Temperature of rotator was too high.

04:37 UT: Classification of AT2016jel. Seeing ~1.3". Wind ~ 11 m/s. Normal Ia, a few days pre max. z~0.06

04:49 UT: Follow up of SN2016jca. 2x2700, gr13. Seeing ~1.3"

06:38 UT: Standard star gd71. gr13+1 only.

06:40 UT: Classification of AT2016jbu. Seeing ~1.1". Likely LBV, or Type II. z~0.01

07:05 UT: R_dit of AT2016jbu.

07:54 UT: Classification of OGLE17abx. Wind ~10m/s. Seeing ~0.9". Type II a few weeks past max. z~0.07.

08:26 UT: Classification of AT2016jgb. Transient missing on 2min acquisition image.

08:37 UT: End of night

08:38 UT: Morning skyflats for R passband.